3. Tzatziki contains 2.4 grams of protein, 0.6 grams of fat, and 1.5 grams of carbohydrate in each
serving. Each gram of fat that is metabolized in the body produces 9.0 Cal of energy, while each gram of
protein or carbohydrate that is metabolized produces 4.0 Cal of energy.
(a) How many Calories (to the nearest Calorie) are provided by one serving of tzatziki?
(b) How many kilojoules (to the nearest kJ) of energy are provided by one serving of tzatziki?
(c) What percentage of the energy in this food is provided by the fat? (to 2 s.f.)
1
Expert's answer
2020-09-29T06:48:33-0400

1 gm of protein produces 4 calories of energy so 2.4 gm protein = 9.6 calories of energy.

0.6 gm of fat produces 9x0.6 = 5.4 calories energy.

1.5 gm carbohydrate produces 4x1.5 = 6 calories of energy.

(a) so Calories provided by one serving of tzatziki is 9.6+5.4+6 = 21 calories.

(b) same above calories 21 calories = 4.184x21x10^(-3) KJ = 8.8 * 10^(-2) KJ.

(c) percentage of fat = (5.4/21)x100 = 25.7 %